Retail price of fuel is uniformly distributed between $2.00 and $3.50
A) Calculate the retail average price of fuel?
B) what is the probability that the price of fuel is less than $3.00
C) what is the probability that price of fuel is less than $2.00
1
Expert's answer
2020-05-03T16:58:36-0400

A) The retail average price of fuel is:

"P = (2 + 3.5)\/2 = 2.75."

B) The probability that the price of fuel is less than $3.00 is:

"\frac{3 - 2} {3.5 - 2} = 2/3."

C) The probability that price of fuel is less than $2.00 is:

"\frac{2-2} {3.5-2} = 0."